1.    How many atoms of carbon are there in 0.40 mol of procaine, C13H20N2O2, a “pain killer” used by dentists?


1
Expert's answer
2021-09-16T02:04:28-0400

1 mole of procaine contains "6.022\times 10^{23}molecules"

0.4 mole of procaine contains "0.4\times6.022\times 10^{23} molecules"

Therefore, 0.4 mol of procaine contains "\times 6.022\times 10^{23}\times 13" atoms of carbon "=3.1\times 10^{24}"

atoms of carbon.